OpenStudy (anonymous):

What is the unit fraction of 9/10?

OpenStudy (anonymous):

what is unit fraction?

OpenStudy (anonymous):

A egyptian way of writing a fraction the numerator has to be 1

OpenStudy (anonymous):

and the denominator cant repeat

OpenStudy (anonymous):

for ex: 4/5 =1/2 + 1/5 + 1/10

OpenStudy (anonymous):

i have only 1/10 + 1/2+ 1/5 so far

OpenStudy (anonymous):

interesting :D

OpenStudy (anonymous):

...

OpenStudy (anonymous):

do you need to guess :D or there are methodology how to make them?

OpenStudy (anonymous):

umm i think guess...?

OpenStudy (anonymous):

idk.

OpenStudy (anonymous):

there isnt really a method

OpenStudy (anonymous):

1/2+1/3+1/15=9/10 http://www.maths.surrey.ac.uk/hosted-sites/R.Knott/Fractions/egyptian.html i think there should be since it's calculated

OpenStudy (anonymous):

1/2 + 1/4 + 1/8 + 1/40

OpenStudy (anonymous):

oh?

OpenStudy (anonymous):

using greedy algorhitm

OpenStudy (anonymous):

..

OpenStudy (anonymous):

Thanks guys

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

95/14541 = 1/154+1/25161+1/3756225304+1/35273071332268507000 :D

OpenStudy (anonymous):

uh

OpenStudy (anonymous):

im kinda confused?

OpenStudy (anonymous):

D:<

OpenStudy (anonymous):

there is even egyptian fraction for a \(\pi\)!

OpenStudy (anonymous):

omg..

OpenStudy (anonymous):

wait how is 1/2 1/3 1/15 da answer

OpenStudy (anonymous):

8, 61, 5020, 128541455, 162924332716605980, ...

OpenStudy (anonymous):

coould u turn it into equalivent fracs

OpenStudy (anonymous):

\[\frac{1}{2}+\frac{1}{3}+\frac{1}{15}=\frac{15}{30}+\frac{10}{30}+\frac{2}{30}=\frac{15+10+2}{30}=\frac{27}{30}=\frac{9}{10}\]

OpenStudy (anonymous):

can't you do it yourself?

OpenStudy (anonymous):

holy

OpenStudy (anonymous):

OHHHHHHHHHH i get it

OpenStudy (anonymous):

THANKS SO MUCHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHH!!!!!!!!!

OpenStudy (anonymous):

thank you for showing my egyptian fractions :D

OpenStudy (anonymous):

yup :P